Ca2+ standard/mL Absorbance Ca2+ standard/mL Absorbance
0.000 0.151 8.00 0.388
1.00 0.185 10.00 0.445
3.00 0.247 15.00 0.572
5.00 0.300 20.00 0.723

In determining the Ca content in a breakfast cereal, 0.5216 g of crushed cereal was ashed in a crucible at 600 degrees celcius in air for 2 hours. The residue was dissolved in 6 mol/L HCl, quantitatively transferred to a volumetric flask, and diluted to 100.0 mL. Then, 5.00-mL aliquots were transferred to 50-mL volumetric flasks. Each was treated with Ca2+ standard (containing 20.0 micrograms/mL), diluted to volume with H2O, and analysed by flame atomic absorption. The resultss obtained are tabulated below. Using these results, find the weight percentage of Ca in the cereal and its 95% confidence interval.
Answer #1

Plot a Absorbance vs standard added graph and calculate the value of absorbance at y= 0

2 L 2.1 CD 焦慮》 2i0 1.1 standard added y= 0.0282x+ 0.1579

value of x at y=0 is : x = -5.599

concentration of unknown = - (-5.599) * (volume of flask/volume of unknown)

                                          = 5.599 * (50/5)

                                         = 55.99 mL

concentration of Ca^2+ in the 5mL flak = 55.99 *20microgram/mL *10

concentration of ca^2+ in 100mL flask = 224000microgram = 0.224 gm

% Ca= 0.224/0.5216 *100% = 42.9%